OpenStudy (anonymous):

(a)/(-3) ÷ (5b)/(-3)

sam (.sam.):

\[\frac{a}{-3} \div \frac{5b}{-3} \\ \\ Flip \\ \\ \frac{a}{-3} \times \frac{-3}{5b}\] Cross multiply now

OpenStudy (anonymous):

Is it a over 5b

sam (.sam.):

Good!